# On uniqueness in structured model learning

### v2

@@ B.2 Nonlinear case
Boundedness of weakly convergent sequences (see e.g. [8, Proposition 3.5 (iii)]) and $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ together with the assumptions on $\mathcal{B}_{0}$ ensure the existence of $c_{\varphi},c_{v}>0$ such that both $\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{\mathcal{C}(0,T;H)})\leq\mathcal{B}_{0}(c_{\varphi},c_{v})$ and
@@ Appendix C Existence of minimizers
By Assumption 2, iv), the map $\gamma$ is continuous. Together with separability of the spaces $V,B$ and $p\geq s$ we derive by [44, Theorem 1.43] that $\gamma$ induces a well defined Nemytskii operator $\gamma:L^{p}(0,T;V)\to L^{s}(0,T;B)=\mathcal{B}$ which is continuous. Employing $\mathcal{V}\hookrightarrow L^{p}(0,T;V)$ and linearity of $\gamma$ concludes the proof. ∎
@@
Note that problem ($\mathcal{P}^{\prime}$) is in canonical form as the sum of a data-fidelity term and a regularization functional where $G$, given in (60), is the forward operator and $(0,0,0,y)\in\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$ the measured data. We prove that problem ($\mathcal{P}^{\prime}$) admits a solution in $\mathbf{D}(G)$. If the forward operator $G$ is weakly closed then problem ($\mathcal{P}^{\prime}$) admits a minimizer due to the direct method (see e.g. [51, Chapter 3]) and Assumption 2, vi). The idea is to choose a minimizing sequence, which certainly, for indices large enough is bounded by coercivity of the regularizer, the norm in $H$ and the discrepancy term (together with boundedness of the trace map), thus, attaining a weakly convergent subsequence. Employing weak closedness of $G$, weak lower semicontinuity of the norms, the regularizing term and the discrepancy term (due to Assumption 2, i) and Lemma 43) we derive that the limit of this subsequence is a solution of the minimization problem ($\mathcal{P}^{\prime}$).
@@ Appendix D Proof of Proposition 1
for a sufficiently large interval $U$ and regularization parameters $\nu^{m}>0$, is coercive and weakly lower semicontinuous (which follows by Proposition 19). Note that $N=1$. Furthermore, the choice of the image space as $W=L^{2}(\Omega)$ is justified since $\partial_{t}u-\varphi\cdot\nabla u\in L^{2}(\Omega)$ for $\varphi\in H^{1}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ and $u\in\mathcal{V}=W^{1,2,2,}(0,T;V)$ as $\partial_{t}u\in L^{2}(\Omega)$ and $\nabla u\in L^{2}(\Omega)$ by $u\in\mathcal{V}=W^{1,2,2,}(0,T;V)$. The inclusion for parameterized nonlinearities $f_{\theta}(u)\in L^{2}(\Omega)$ is fulfilled by the following considerations.
@@ Appendix E Example on choice of regularization
In this section we showcase the necessity of choosing a $W^{1,\infty}$-type norm in accordance with (3) for identifying hidden physics based on an example. For that, we consider a one dimensional time-independent equation on the unit interval with known physical term $F$ that depends on the state and its first spatial derivative. Furthermore, we suppose that the state $u$ is given via the full measurement operator (e.g. equal to the identity with noiseless measurement data) and is approximated by a known sequence of states $(u_{m})_{m}$ in $L^{2}(0,1)$. Note that, although this setup is simpler than the general one considered in this work, the example shows that already in this simplified situation appropriate regularization, as discussed, matters. We suppose that $(-F(u_{m},u_{m}^{\prime}))_{m}$ converges to some $g$ in $L^{2}(0,1)$ (this condition is verified for the explicitly chosen terms below). Furthermore, we consider classes of continuous functions parameterized by sets $\Theta^{m}$ that can approximate more and more complicated functions better for increasing $m\in\mathbb{N}$ (according to Assumption 5, ii) and iv)) such that any constant function can be explicitly represented for sufficiently large $m\in\mathbb{N}$. With this, we provide an example where solutions $(f_{\theta^{m}})_{m}$ with $\theta^{m}\in\Theta^{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m},u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|$ ($\mathcal{P}^{m}_{L^{2}}$)
do not converge to the unique solution $f^{\dagger}\in L^{2}(0,1)$ of
@@
It is important to note that the constraint in ($\mathcal{P}^{\dagger}_{L^{2}}$) is not well-defined for general $f,u,g\in L^{2}(0,1)$ but for the setup discussed below it is. In view of well-definedness of the minimization problem ($\mathcal{P}^{m}_{L^{2}}$) we refer to the considerations below. Here $(\lambda^{m})_{m},(\mu^{m})_{m}\subset\mathbb{R}_{+}$ are suitable regularization parameters, the former monotonically increasing and divergent and the latter a zero sequence. Finally, in contrast to above situation we show that penalizing the gradient term similarly as in (3) allows to identify the unique solution $\tilde{f}^{\dagger}$ of
@@
as limit of solutions $(f_{\tilde{\theta}^{m}})_{m}$ for $\tilde{\theta}^{m}\in\Theta^{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(0,1)}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m},u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|.$ ($\mathcal{P}^{m}_{W^{1,\infty}}$)
Concretely we choose the known physical term $F$ to be given by
@@
$\mathbb{R}^{2}\ni(v,w)\mapsto F(v,w)=\begin{cases}\frac{v}{w}-1,&\text{if}~w\neq 0\\ -1,&\text{otherwise}\end{cases}.$
Furthermore, we suppose that the state is given by $u=0$ on $[0,1]$ and approximated by the sequence of states $u_{m}(x)=x^{m}$ for $m\in\mathbb{N}$. In fact $(u_{m})_{m}$ converges to $u$ in $L^{2}(0,1)$ since $\|u_{m}\|_{L^{2}(0,1)}^{2}=(2m+1)^{-1}$ for $m\in\mathbb{N}$. As a consequence, we derive that for $g_{m}:=-F(u_{m},u_{m}^{\prime})$, i.e. $g_{m}(x)=1-\frac{1}{m}x$ for $x\in[0,1]$ and $m\in\mathbb{N}$, the sequence $(g_{m})_{m}$ converges to $g=1$ in $L^{2}(0,1)$ since $\|1-g_{m}\|_{L^{2}(0,1)}^{2}=(3m^{2})^{-1}$. The regularization parameters are chosen by $\lambda_{m}=\lambda_{0}m^{3/2}$ for $m\in\mathbb{N}$ and fixed $\lambda_{0}>0$ whereas $(\mu_{m})_{m}$ is a zero sequence such that with $\hat{\theta}^{m}\in\Theta^{m}$ parameterizing the constant one function, it holds that $(\mu_{m}\|\hat{\theta}^{m}\|)_{m}$ is a zero sequence. Note that the PDE data term $\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ is even defined for $f\in L^{2}(0,1)$ as follows. For continuous $f\in\mathcal{C}(0,1)$ we derive by integrating via substitution that
$\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}=\frac{1}{m}\int_{0}^{1}x^{\frac{m}{m-1}}\left(f(x)-(1-m^{-1}x^{1/m})\right)^{2}\mathop{}\!\mathrm{d}x.$ (61)
@@
Since $|x^{\frac{m}{m-1}}|\leq 1$ and $|1-m^{-1}x^{1/m}|\leq 1$ for $x\in[0,1]$, it follows that the right hand side of (61) is well-defined also for $f\in L^{2}(0,1)$. For $f\in W^{1,\infty}(0,1)$ well-definedness follows directly due to the embedding $W^{1,\infty}(0,1)\hookrightarrow\mathcal{C}(0,1)$. Furthermore, note that well-definedness of ($\mathcal{P}^{m}_{W^{1,\infty}}$) follows by the considerations in this work, whereas well-definedness of ($\mathcal{P}^{m}_{L^{2}}$) is a consequence of the direct method together with (61) applied to parameterizations instead of $f$ and continuity w.r.t. to the parameterization similar as in Proposition 18.
@@
Now for sufficiently large $m\in\mathbb{N}$ the constant one function is exactly representable via some $f_{\hat{\theta}^{m}}$ with $\hat{\theta}^{m}\in\Theta^{m}$ such that the objective functional of ($\mathcal{P}^{m}_{L^{2}}$) in an optimum can be estimated by
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\lambda_{0}m^{3/2}\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\theta^{m}\|\\ \leq 1+\lambda_{0}m^{3/2}\|1-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\hat{\theta}^{m}\|.$
@@
Due to the chosen setup the right hand side is uniformly bounded for sufficiently large $m\in\mathbb{N}$. Thus, it follows that $m\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ is a zero sequence, which with (61) implies that $x\mapsto x^{\frac{m}{2(m-1)}}(f_{\theta^{m}}(x)-(1-m^{-1}x^{1/m}))$ converges to zero in $L^{2}(0,1)$. Hence, a subsequence (w.l.o.g. the entire sequence) converges a.e. in $[0,1]$ to zero. Since $x\mapsto 1-m^{-1}x^{1/m}$ converges to the constant one uniformly in $[0,1]$ and $x\mapsto x^{\frac{m}{2(m-1)}}$ to $x\mapsto x^{1/2}$ it follows that $f_{\theta^{m}}(x)$ converges to the constant one a.e. in $[0,1]$. However, the constant one function is not the unique solution to ($\mathcal{P}^{\dagger}_{L^{2}}$), which is given by the constant zero function (note that in fact even continuous functions with smaller $L^{2}$-norm than the constant one function fulfilling the constraint of ($\mathcal{P}^{\dagger}_{L^{2}}$) exist such as $x\mapsto 1-x$). As a consequence, the reconstruction of the hidden physical term fails. In fact one can show for the minimization problem similar to ($\mathcal{P}^{m}_{L^{2}}$) considered over general $L^{2}(0,1)$-functions (and without parameter regularization) that by analyzing the first variation of the resulting strictly convex objective functional under (61), the corresponding minimizer $f_{m}$ is given by
$f_{m}(x)=\lambda_{0}\left(\lambda_{0}+m^{-1/2}x^{\frac{m-1}{m}}\right)^{-1}\left(1-\frac{1}{m}x^{1/m}\right)\quad\text{for}~x\in[0,1],$
@@
which can be easily shown to converge to the constant one in $L^{2}(0,1)$. Another interesting point is that the $(f_{m})_{m}$ are continuously differentiable in the open unit interval and $\|\nabla f_{m}\|_{L^{\infty}(0,1)}=\infty$ for $m\in\mathbb{N}$. Furthermore, one can provide the lower bound estimation $|\nabla f_{m}(m^{-3})|\geq m/2$ for sufficiently large $m\in\mathbb{N}$. Thus, also in a purely analytic setup, the reconstruction of the hidden physics fails.
Let us consider next the $W^{1,\infty}$-type regularized problems, where we first address in view of ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) for fixed $c\geq 0$ the problem
$\min_{f\in W^{1,\infty}(0,1)}\|f\|_{L^{2}(0,1)}^{2}\quad\text{s.t.}~f(0)=1~\text{and}~\|\nabla f\|_{L^{\infty}(0,1)}=c.$
Due to the $L^{2}$-term the solution is given by the monotonically decreasing function $f_{c}$ which fulfills $f_{c}(x)=1-cx$ if $cx\leq 1$ and zero otherwise, i.e. $f_{c}$ attains maximally possible negative slope as long as the $L^{2}$-norm is reduced. Now problem ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) is equivalent to $\min_{c\geq 0}\|f_{c}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{c}\|_{L^{\infty}(0,1)}$, since only the functions $(f_{c})_{c}$ come into question for minimization, which by $\|f_{0}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{0}\|_{L^{\infty}(0,1)}=1<\|\nabla f_{c}\|_{L^{\infty}(0,1)}$ for $c>1$ is further equivalent to
$\min_{0\leq c\leq 1}\|f_{c}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{c}\|_{L^{\infty}(0,1)}=\min_{0\leq c\leq 1}(1-c+c^{2}/3)+c=1.$
@@
As a consequence, the unique solution to ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) is given by $\tilde{f}^{\dagger}=f_{0}=1$, the constant one function. We prove that $\tilde{f}^{\dagger}$ is recovered by $(f_{\tilde{\theta}^{m}})_{m}$ solving ($\mathcal{P}^{m}_{W^{1,\infty}}$). For that, we first estimate the objective function of ($\mathcal{P}^{m}_{W^{1,\infty}}$) in the optimum by its value at $f_{\hat{\theta}^{m}}$ representing the constant one function, yielding

### v3

@@ B.2 Nonlinear case
Boundedness of weakly convergent sequences (see e.g. [15, Proposition 3.5 (iii)]) and $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ together with the assumptions on $\mathcal{B}_{0}$ ensure the existence of $c_{\varphi},c_{v}>0$ such that both $\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{\mathcal{C}(0,T;H)})\leq\mathcal{B}_{0}(c_{\varphi},c_{v})$ and
@@ Appendix C Existence of minimizers
By Assumption 2, iv), the map $\gamma$ is continuous. Together with separability of the spaces $V,B$ and $p\geq s$ we derive by [88, Theorem 1.43] that $\gamma$ induces a well defined Nemytskii operator $\gamma:L^{p}(0,T;V)\to L^{s}(0,T;B)=\mathcal{B}$ which is continuous. Employing $\mathcal{V}\hookrightarrow L^{p}(0,T;V)$ and linearity of $\gamma$ concludes the proof. ∎
@@
Note that problem ($\mathcal{P}^{\prime}$) is in canonical form as the sum of a data-fidelity term and a regularization functional where $G$, given in (61), is the forward operator and $(0,0,0,y)\in\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$ the measured data. We prove that problem ($\mathcal{P}^{\prime}$) admits a solution in $\mathbf{D}(G)$. If the forward operator $G$ is weakly closed then problem ($\mathcal{P}^{\prime}$) admits a minimizer due to the direct method (see e.g. [100, Chapter 3]) and Assumption 2, vi). The idea is to choose a minimizing sequence, which certainly, for indices large enough is bounded by coercivity of the regularizer, the norm in $H$ and the discrepancy term (together with boundedness of the trace map), thus, attaining a weakly convergent subsequence. Employing weak closedness of $G$, weak lower semicontinuity of the norms, the regularizing term and the discrepancy term (due to Assumption 2, i) and Lemma 44) we derive that the limit of this subsequence is a solution of the minimization problem ($\mathcal{P}^{\prime}$).
@@ Appendix D Proof of Proposition 1
for a sufficiently large interval $U$ and regularization parameters $\nu^{m}>0$, is coercive and weakly lower semicontinuous (which follows by Proposition 20). Note that $N=1$. Furthermore, the choice of the image space as $W=L^{2}(\Omega)$ is justified since $\partial_{t}u-\varphi\cdot\nabla u\in L^{2}(\Omega)$ for $\varphi\in H^{1}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ and $u\in\mathcal{V}=W^{1,2,2}(0,T;V)$ as $\partial_{t}u\in L^{2}(\Omega)$ and $\nabla u\in L^{2}(\Omega)$ by $u\in\mathcal{V}=W^{1,2,2}(0,T;V)$. The inclusion for parameterized nonlinearities $f_{\theta}(u)\in L^{2}(\Omega)$ is fulfilled by the following considerations.
@@ Appendix E Example on choice of regularization
In this section, based on a simple example, we show that choosing a $W^{1,\infty}$-type norm in accordance with (3), as opposed to an $L^{p}$-type norm, can indeed be necessary in general for the identification of a hidden physics component. For that, we consider a one dimensional time-independent equation on the unit interval with known physical term $F$ that depends on the first spatial derivative of the state. Furthermore, we suppose that the state $u$ is given via the full measurement operator (e.g. equal to the identity with noiseless measurement data) and is approximated by a known sequence of states $(u_{m})_{m}$. Note that, although this setup is simpler than the general one considered in this work, the example shows that already in this simplified situation appropriate regularization, as discussed, matters. Considering classes of continuous functions parameterized by sets $\Theta^{m}$ that can approximate more and more complicated functions better for increasing $m\in\mathbb{N}$ (according to Assumption 5, ii) and iv)), we provide an example where solutions $(f_{\theta^{m}})_{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|$ ($\mathcal{P}^{m}_{L^{2}}$)
with $\theta^{m}\in\Theta^{m}$, do not converge to the unique solution $f^{\dagger}\in L^{2}(0,1)$ of
@@
It is important to note that the constraint in ($\mathcal{P}^{\dagger}_{L^{2}}$) is not well-defined for general $f,u,g\in L^{2}(0,1)$ but for the setup discussed below it is. In view of well-definedness of the minimization problem ($\mathcal{P}^{m}_{L^{2}}$) we refer to the considerations below. Here $(\lambda^{m})_{m},(\mu^{m})_{m}\subset\mathbb{R}_{+}$ are suitable regularization parameters, the former monotonically increasing and divergent, and the latter a zero sequence. Finally, in contrast to above situation we show that penalizing the gradient term similarly as in (3) allows to identify the unique solution $f^{\dagger}$ of
@@
as limit of solutions $(f_{\theta^{m}})_{m}$ for $\theta^{m}\in\Theta^{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(0,1)}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|.$ ($\mathcal{P}^{m}_{W^{1,\infty}}$)
Concretely, we choose the known physical term $F$ by
@@
$\mathbb{R}\ni v\mapsto F(v)=\begin{cases}\left(1-4(v-1)^{2}\right)^{1/2},&\text{if}~1/2\leq v\leq 3/2,\\ 0,&\text{otherwise.}\end{cases}$
Furthermore, we suppose that the state is given by the identity map $u(x)=x$ for $x\in[0,1]$ and is approximated by the sequence of states $(u_{m})_{m}$ given for $m\in\mathbb{N}$ by
$u_{m}(x)=x+\frac{1}{4\pi m}\sin(2\pi mx)\quad\text{for}~x\in[0,1].$
The sequence $(u_{m})_{m}$ converges to $u$ in $L^{\infty}(0,1)$ since $\|u_{m}-u\|_{L^{\infty}(0,1)}=(4\pi m)^{-1}$ for $m\in\mathbb{N}$ (in fact, even weakly in $H^{1}(0,1)$). Note that $u_{m}:[0,1]\to[0,1]$ is bijective and even diffeomorphic since $u_{m}^{\prime}(x)=1+\cos(2\pi mx)/2\in[1/2,3/2]$ for all $x\in[0,1]$ and $m\in\mathbb{N}$. The sequence $(g_{m})_{m}$ with $g_{m}:=-F(u_{m}^{\prime})$, i.e., for $m\in\mathbb{N}$
$g_{m}(x)=\sin(2\pi mx)\quad\text{for}~x\in[0,1],$
converges to $g=0$ weakly in $L^{2}(0,1)$. The regularization parameters are chosen by $\lambda^{m}=\lambda_{0}m^{1/2}$ for $m\in\mathbb{N}$ and fixed $\lambda_{0}>0$ whereas $(\mu^{m})_{m}$ is a zero sequence satisfying the following conditions. We assume that the zero function can be parameterized for any $m\in\mathbb{N}$ with suitable parameters $\hat{\theta}^{m}\in\Theta^{m}$. Furthermore, suppose that the periodic functions $h_{m}:=g_{m}\circ u_{m}^{-1}$ (which have period $1/m$) can be parameterized with parameters $\tilde{\theta}^{m}\in\Theta^{m}$ for $m\in\mathbb{N}$. Then we choose $(\mu^{m})_{m}$ such that both $(\mu^{m}\|\hat{\theta}^{m}\|)_{m}$ and $(\mu^{m}\|\tilde{\theta}^{m}\|)_{m}$ are zero sequences. Note that the PDE data term $\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ is well-defined for any $f\in L^{2}(0,1)$. Indeed, since $u_{m}$ is a diffeomorphism of the interval $[0,1]$ onto itself, a change of variables yields
$\|f(u_{m})\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}|f(x)|^{2}(u_{m}^{\prime}(u_{m}^{-1}(x)))^{-1}\mathop{}\!\mathrm{d}x.$ (62)
Using that $1/2\leq u_{m}^{\prime}(x)\leq 3/2$ for $x\in[0,1]$ and $m\in\mathbb{N}$ implies
$\frac{2}{3}\|f\|_{L^{2}(0,1)}^{2}\leq\|f(u_{m})\|_{L^{2}(0,1)}^{2}\leq 2\|f\|_{L^{2}(0,1)}^{2}.$ (63)
Note that by similar arguments we derive for $h_{m}=g_{m}\circ u_{m}^{-1}$ that
$\|h_{m}\|_{L^{2}(0,1)}^{2}\leq\frac{3}{2}\|g_{m}\|_{L^{2}(0,1)}^{2}=3/4$ (64)
@@
using that $\|g_{m}\|_{L^{2}(0,1)}^{2}=1/2$ for $m\in\mathbb{N}$. For $f\in W^{1,\infty}(0,1)$ well-definedness follows from the embedding $W^{1,\infty}(0,1)\hookrightarrow\mathcal{C}(0,1)$. Furthermore, note that well-definedness of ($\mathcal{P}^{m}_{W^{1,\infty}}$) follows by the considerations in this work, whereas well-definedness of ($\mathcal{P}^{m}_{L^{2}}$) is a consequence of the direct method together with (62) applied to parameterizations instead of $f$ and continuity with respect to the parameterization similar as in Proposition 19.
@@
We now start by considering the $L^{2}$-regularized problem. For $m\in\mathbb{N}$ the function $h_{m}=g_{m}\circ u_{m}^{-1}$ is representable by $f_{\tilde{\theta}^{m}}$ with $\tilde{\theta}^{m}\in\Theta^{m}$ such that the objective functional of ($\mathcal{P}^{m}_{L^{2}}$) in an optimum can be estimated by
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\lambda_{0}m^{1/2}\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\theta^{m}\|\\ \leq\|f_{\tilde{\theta}^{m}}\|_{L^{2}(0,1)}^{2}+\lambda_{0}m^{1/2}\|\underbrace{f_{\tilde{\theta}^{m}}(u_{m})-g_{m}}_{=0}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\tilde{\theta}^{m}\|.$
Due to the choice of $(\mu^{m})_{m}$ and (64) the right hand side is uniformly bounded for sufficiently large $m\in\mathbb{N}$. Thus, it follows that $\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ converges to zero as $m\to\infty$. Hence, there exists a constant $c>0$ such that by the reverse triangle inequality for sufficiently large $m\in\mathbb{N}$
$\|f_{\theta^{m}}(u_{m})\|_{L^{2}(0,1)}\geq\underbrace{\|g_{m}\|_{L^{2}(0,1)}}_{=2^{-1/2}}-\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}\geq c>0.$ (65)
@@
Thus, the sequence $(f_{\theta^{m}})_{m}$ cannot converge to zero in $L^{2}(0,1)$ as (63) would immediately lead to a contradiction to (65). However, the constant zero function is the unique solution to ($\mathcal{P}^{\dagger}_{L^{2}}$) (of course up to representatives in the Lebesgue sense). As a consequence, the reconstruction of the hidden physical term fails. In fact, one can show for the minimization problem similar to ($\mathcal{P}^{m}_{L^{2}}$) considered over general $L^{2}(0,1)$-functions (and without parameter regularization) that by analyzing the first variation of the resulting strictly convex objective functional under (62), the corresponding minimizer $f_{m}$ is given by
$f_{m}(x)=\lambda_{0}\left(\lambda_{0}+m^{-1/2}u_{m}^{\prime}(u_{m}^{-1}(x))\right)^{-1}g_{m}(u_{m}^{-1}(x))\quad\text{for}~x\in[0,1],$
@@
whose $L^{2}(0,1)$-norm can be shown to be uniformly bounded from below. Another interesting point is that the $(f_{m})_{m}$ are continuously differentiable in the open unit interval and $\lim_{m\to\infty}\|\nabla f_{m}\|_{L^{\infty}(0,1)}=\infty$. Thus, also in a purely analytic setup, the reconstruction of the hidden physics fails.
@@
Let us consider next the $W^{1,\infty}$-type regularized problems. Problem ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) attains the unique solution $f^{\dagger}=0$. We argue that $f^{\dagger}$ is recovered by $(f_{\theta^{m}})_{m}$ solving ($\mathcal{P}^{m}_{W^{1,\infty}}$). For that, we first estimate the objective function of ($\mathcal{P}^{m}_{W^{1,\infty}}$) in the optimum by its value at $f_{\hat{\theta}^{m}}$ representing the constant zero function, yielding
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}+\lambda_{0}m^{1/2}\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\theta^{m}\|\\ \leq\lambda_{0}m^{1/2}\|g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\hat{\theta}^{m}\|=\lambda_{0}m^{1/2}/2+\mu^{m}\|\hat{\theta}^{m}\|.$ (66)
We derive by (66) that
$\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq\lambda_{0}m^{1/2}/2+\mu^{m}\|\hat{\theta}^{m}\|,$
which implies the existence of a constant $c>0$ such that $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq cm^{1/2}$ for $m\in\mathbb{N}$ since $(\mu^{m}\|\hat{\theta}^{m}\|)_{m}$ is a zero sequence. Under this constraint on the gradient, the best approximation of the data term is bounded from below by
$\inf_{\begin{subarray}{c}\theta\in\Theta^{m},\\ \|\nabla f_{\theta}\|_{L^{\infty}(0,1)}\leq c\sqrt{m}\end{subarray}}\|f_{\theta}(u_{m})-g_{m}\|_{L^{2}(0,1)}\geq\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq c\sqrt{m}\end{subarray}}\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}.$ (67)
We analyze the right hand side in more detail. By the mean value theorem it holds true for $f\in W^{1,\infty}(0,1)$ that
$\|f(u)-f(u_{m})\|_{L^{2}(0,1)}\leq\|\nabla f\|_{L^{\infty}(0,1)}\|u-u_{m}\|_{L^{2}(0,1)}=\tilde{c}\|\nabla f\|_{L^{\infty}(0,1)}m^{-1}$
with $\tilde{c}=32^{-1/2}\pi^{-1}$, such that (67) can be estimated from below by
$\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq cm^{1/2}\end{subarray}}\|f-g_{m}\|_{L^{2}(0,1)}-\tilde{c}cm^{-1/2}.$ (68)
Using a scaling argument, the value of the approximation problem in (68) equals
$cm^{1/2}\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq 1\end{subarray}}\|f-c^{-1}m^{-1/2}g_{m}\|_{L^{2}(0,1)}.$ (69)
Due to [18, Theorem 1.1] this problem attains a unique solution $\hat{f}$ which satisfies
$\hat{f}(x)$ $=\max_{y\in\partial A^{+}}\hat{f}(y)-|x-y|\qquad\text{for all}~x\in A^{+},$
$\hat{f}(x)$ $=\min_{y\in\partial A^{-}}\hat{f}(y)+|x-y|\qquad\text{for all}~x\in A^{-},$
where $A^{+}=\text{supp}(\max(\hat{f}-c^{-1}m^{-1/2}g_{m},0))$ and $A^{-}=\text{supp}(\min(\hat{f}-c^{-1}m^{-1/2}g_{m},0))$ denote the respective supports, and $\partial A^{+},\partial A^{-}$ the corresponding boundaries. With this, and the fact that $c^{-1}m^{-1/2}g_{m}$ attains $m$ full periods in $[0,1]$, it follows by a symmetry argument that $\hat{f}$ is given by the saw-tooth function with unit slope and same period and sign as $g_{m}$. As a consequence, the term in (69) equals
$cm^{1/2}\left(4m\int_{0}^{1/4m}(c^{-1}m^{-1/2}g_{m}(x)-x)^{2}\mathop{}\!\mathrm{d}x\right)^{1/2}=\left(\frac{1}{2}-\frac{2c}{\pi^{2}m^{1/2}}+\frac{c^{2}}{48m}\right)^{1/2}.$
Combining these arguments, finally, yields by (66) that